For two vectors, what angle between them will lead to smallest magnitudes when added?
1
Expert's answer
2020-07-21T12:39:35-0400

Let's consider the sum of two vectors "\mathbf{a}" and "\mathbf{b}" with the angle "\alpha" between them:



Here "\mathbf{c} = \mathbf{a} +\mathbf{b}". Using the law of cosines, we can find the length of the "\mathbf{c}":


"c^2 = a^2 +b^2-2ab\cos(180\degree-\alpha)"

Using the trig identity:


"\cos(180\degree-\alpha) = -\cos\alpha"

obtain:


"c^2 = a^2 +b^2+2ab\cos\alpha"

In order to minimize the "c", the "\alpha" should be chosen equal to "180\degree". Then "\cos\alpha = -1" and


"c^2 = a^2 +b^2-2ab"

Answer. The angle 180 between two vectors will lead to smallest magnitudes when added.